OpenStudy (anonymous):

how the gdhajgdkgdbsjkkjabchbsxhjzsVxJZBK do i find the vertex of a parabola? do i just solve?

OpenStudy (anonymous):

...

OpenStudy (accessdenied):

in standard form (ax^2 + bx + c), the vertex is at x= -b/2a in vertex form (a(x - h)^2 + k), the vertex is at x = h

OpenStudy (anonymous):

wtf is h and k?

OpenStudy (accessdenied):

coordinates for the vertex of some arbitrary function in vertex form?

OpenStudy (anonymous):

k